OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below:

OpenStudy (anonymous):

What's the example?

OpenStudy (anonymous):

my bad it is Solve 4y - y + 11 = 38

OpenStudy (anonymous):

First subtract the y. 4y - y + 11 = 38. 3y + 11 = 38. Then subtract the 11 from the other side to isolate the y. 3y = 27. Divide both sides by 3 to isolate the y. y = 27/3. y = 9.

OpenStudy (anonymous):

think you so much

OpenStudy (anonymous):

can you help me with one more

OpenStudy (anonymous):

x: 1/4 (2x − 14) = 4

OpenStudy (anonymous):

Solving for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/4 (2x-14) = 4. The best way to start would be to get rid of the fraction, so multiply both sides by 4, and you get 2x-14 = 16. Do the same thing we did in the other problem, and isolate the variable by adding 14 to both sides. 2x = 30. x = 15.

OpenStudy (anonymous):

thanks so much